# Nemotron-CrossThink: Scaling Self-Learning beyond Math Reasoning

### v2

@@ Abstract
Large Language Models (llms) have shown strong reasoning capabilities, particularly when enhanced through Reinforcement Learning (rl). While prior work has successfully applied rl to mathematical reasoning—where rules and correctness are well-defined—generalizing these methods to broader reasoning domains remains challenging due to limited data, the lack of verifiable reward structures, and diverse task requirements. In this work, we propose Nemotron-CrossThink, a framework that systematically incorporates multi-domain corpora, including both synthetic and real-world question-answer pairs, into rl training to improve generalization across diverse reasoning tasks. Nemotron-CrossThink addresses key challenges by (1) incorporating data from varied sources spanning STEM, humanities, social sciences, etc.; (2) applying structured templates (e.g., multiple-choice and open-ended) to control answer-space complexity; (3) filtering for verifiable answers; and (4) optimizing data blending strategies that utilizes data from multiple sources effectively. Our approach enables scalable and verifiable reward modeling beyond mathematics and demonstrates improved accuracies on both math (math-500: +30.1%, amc23: +27.5%) and non-math reasoning benchmarks (mmlu-pro: +12.8%, gpqa-diamond: +11.3%, agieval: +15.1%, supergpqa: +3.8%). Moreover, Nemotron-CrossThink exhibits significantly improved response efficiency—using 28% fewer tokens for correct answers—highlighting more focused and effective reasoning. Through Nemotron-CrossThink, we demonstrate that integrating multi-domain, multi-format data in rl leads to more accurate, efficient, and generalizable llms.
@@ 1 Introduction
Large Language Models (llms) have demonstrated remarkable reasoning capabilities across a wide range of tasks, with Reinforcement Learning (rl) playing a crucial role in refining their deep thinking abilities (Hu et al., 2025; Aggarwal & Welleck, 2025; Luo et al., 2025; DeepSeek-AI, 2025; Qin et al., 2024; Huang et al., 2025; Team, 2025b). Recent advances in rl have been particularly successful in mathematical reasoning and coding, where well-defined rules and verifiable correctness criteria enable effective reward modeling. However, extending these techniques to broader reasoning domains presents significant challenges, such as—including limited training data for rl due to the difficulty of defining verifiable rewards, and ensuring generalization across diverse tasks.
Recent work (Hu et al., 2025; Luo et al., 2025; Cui et al., 2025) has shown a way to diversify rl training corpora by collecting datasets from multiple sources. However, they do not evaluate the relative importance of each source for downstream tasks, nor do they explore optimal data-blending strategies to maximize performance gains. Furthermore, prior research has largely focused on mathematical reasoning, overlooking the impact of incorporating non-math reasoning domains in rl-based learning for generalization in out-of-distribution domains. A major challenge in applying rl to general-purpose reasoning tasks lies in designing a verifiable reward model for diverse answer spaces, as unlike mathematical reasoning—where correctness can be objectively verified—other reasoning tasks lack deterministic solutions. Moreover, reasoning process varies across domains and question types. For instance, mathematical problem-solving follows a rule-based, structured, and symbolic approach (Dehaene, 2011), whereas reasoning in fields such as law, physics, social sciences, and history often relies on narrative structures, contextual knowledge, and heuristic search strategies. Additionally, different question formats require distinct cognitive approaches — open-ended questions demand the generation of novel responses from scratch, while multiple-choice (mcq) questions can often be solved more efficiently by evaluating the given options and selecting the most appropriate answer. Incorporating a diverse range of reasoning domains and question types into rl-based self-learning can enhance the broad reasoning capabilities of llms by exposing them to varied cognitive strategies and knowledge structures.
In this work, we propose Nemotron-CrossThink, a systematic way to incorporate multi-domain corpora for rl training that results in better generalization across a wide variety of tasks. As demonstrated in Figure 2, Nemotron-CrossThink comprises of phases that—(a) curate data from diverse sources, including synthetic data from raw web texts (CommonCrawl) and open-source question-answer pairs, spanning STEM, humanities, law, and social sciences (b) apply templates (mcq/Open-Ended) to limit the answer-space for synthetically generated data (c) filter out samples that are infeasible for verifiable rewards (d) prepare blending recipes to combine different sources of data efficiently and finally (e) employ self-learning with rl to refine reasoning capabilities in diverse domains.
@@
Nemotron-CrossThink demonstrates that integrating multi-domain data with different questions formats for rl significantly enhances reasoning ability of llms across diverse reasoning tasks. Notably, models trained with Nemotron-CrossThink not only achieve higher accuracy but also exhibit dynamic response strategies—generating concise answers for general-purpose questions and more detailed responses for math problems—thereby reducing inference cost while preserving task-specific rigor. In addition, Nemotron-CrossThink addresses the challenge of designing verifiable reward for non-deterministic domains by employing different templates on the curated data to limit the nuances in the answer space diversity. This enables scalable, verifiable reward modeling for general purpose reasoning tasks, ensuring that rl-trained models generalize effectively across diverse benchmarks. Furthermore, Nemotron-CrossThink explores a simple yet effective filtering approach to rank general purpose reasoning data based on complexity and shows that training with harder samples further amplifies the impact of rl across all domains.
In summary, our key contributions are as follows:
@@
  • We introduce Nemotron-CrossThink, a novel framework for incorporating multi-domain corpora into rl training, enhancing the generalization of llms across diverse reasoning tasks with substantial gains across both math (math-500: +30.1%, amc23: +27.5%) and non-math (mmlu-pro: +12.8%, gpqa-diamond: +11.3%, agieval: +15.1%, and supergpqa: +3.8%) benchmarks.
@@
  • We demonstrate that applying question/answer templates to constrain output diversity leads to more stable reward modeling. Specifically, using a unified open-ended question format improves performance by 1.21% on average over mixed-format questions, while short-form answer templates outperform long-form ones by 1.20%.
@@
  • We explore optimal data-blending strategies to balance multi-domain corpora and show that only math data is not enough. Blending multi-domain data boosts average reasoning accuracy by up to 1.61% over math-only training and improves response efficiency by reducing token usage by 28%.
@@
In this paper, we evaluate Nemotron-CrossThink across three dimensions: (1) the effectiveness of different data blending strategies in self-learning (2) whether the blending impact amplifies by filtering and training with more complex data samples (3) the influence of question and answer templates in the downstream performance. Applying Nemotron-CrossThink on different data blends yields substantial improvement over base model, ranging from 8.55% to 13.36% improvement on average across seven diverse general-purpose reasoning and mathematical benchmarks. The most effective blend—constructed using a 2:1 ratio of general-purpose reasoning to math data—achieves the highest average accuracy, improving over the baseline by 13.36% (Figure 1). This underscores the effectiveness of conducting self-learning with combination of data from multiple reasoning domains to enable broader generalization. Our filtering experiment with Qwen-2.5-32B shows a consistent trend, indicating that larger models can further amplify these gains with more complex samples in the data blend (2.15% average improvement), exceeding the improvements observed in the 7B setting. Additionally, our controlled template studies reveal that data formatting decisions play a critical role in model performance. Overall, these findings illustrate that thoughtful choices in data blending, scaling, formatting, and filtering are critical to the success of reinforcement learning with language models. We hope that Nemotron-CrossThink serves as a practical and extensible framework for leveraging multi-domain data to train more capable, reliable, and generalizable models under the rl paradigm.
@@ 2 Nemotron-CrossThink: Scaling Self-Learning Beyond Math
Figure 2: Nemotron-CrossThink. We (a) curate QA pairs from from synthetic (Common Crawl) and open-source datasets, categorized into general-purpose reasoning ($\mathcal{D}_{gpr}$) and mathematical reasoning ($\mathcal{D}_{mr}$); (b) apply structured templates to convert data into multiple-choice (mcq) and open-ended formats, promoting diverse reasoning trajectories; (c) filter out unverifiable or ill-formatted responses; (d) train an RL policy using Group Relative Policy Optimization (grpo). The final reward is used to update the policy, iteratively improving the model’s reasoning capabilities across diverse domains.
In this work, we investigate reasoning domains beyond mathematics and analyze the impact of rl on llms trained with datasets from diverse domains and question formats. A core pre-requisite for effective self-learning is access to high-quality, diverse, and reward-compatible training data (Xie et al., 2025b; Hu et al., 2025). While mathematical reasoning has benefited from clean and verifiable datasets, extending rl to general-purpose reasoning domains remains underexplored due to the lack of structured, high-quality supervision. To address this, we explore methods for leveraging web documents and open-source QA benchmarks to collect general-purpose reasoning data. Incorporating a mix of structured and unstructured domains introduces a wide range of cognitive patterns and task-specific reasoning strategies which will further improve generalization. However, it introduces noise and ambiguity—particularly in open-ended formats—making it difficult to apply rule-based reward modeling reliably. To mitigate this, we apply task-specific templates to unify question and answer formats, limiting answer space variability and enabling simple but effective verifiable reward signals. Next, we apply a lightweight data filtering strategy to discard examples that are infeasible to verify—such as open-ended answers exceeding a certain length or mcqs with misaligned options—ensuring stable and interpretable rl training. Finally, we explore optimal data blending strategies that combine information across diverse domains and tasks. This allows us to investigate how the inclusion of general-purpose reasoning data complements mathematical reasoning, ultimately leading to broader and more adaptive generalization in llms.
@@
We start with carefully curating datasets from multiple sources to ensure diversity in the training data. Our training dataset $\mathcal{D}$ comprises two sources:
@@
Here, $\mathcal{D}_{syn}\rightarrow$ synthetically generated data from Common Crawl (CC) and $\mathcal{D}_{os}\rightarrow$ publicly available open-source QA datasets. Each sources of data further consists of question answer pairs related to general purpose reasoning and mathematics:
$\mathcal{D}_{syn}\rightarrow\mathcal{D}_{syn\_gpr}\cup\mathcal{D}_{syn\_mr};\quad\mathcal{D}_{os}\rightarrow\mathcal{D}_{os\_gpr}\cup\mathcal{D}_{os\_mr}$
@@
Table 1: Training data distribution by source and type. oe=Open-Ended; gpr =General-Purpose Reasoning; mr =Math Reasoning
  • General Purpose Reasoning, $\mathcal{D}_{gpr}$: We collect open source QA datasets ($\mathcal{D}_{os\_gpr}$)—Natural Reasoning (Yuan et al., 2025) and mmlu [Train] (Hendrycks et al., 2021a) that span multiple domains, including STEM fields (e.g., Physics, Computer Science), Economics, Social Sciences, and more. To enhance diversity, we further synthesize QA pairs from CC documents using the wide range of domains in mmlu as our seed domain. We denote this dataset as Syn-qa ($\mathcal{D}_{syn\_gpr}$).
  • Mathematical Reasoning, $\mathcal{D}_{mr}$: As mathematical questions inherently require Chain-of-Thought derivations which emphasizes the llm to think, we incorporate math reasoning corpus to our training data. We combine open-source mathematical reasoning datasets ($\mathcal{D}_{os\_mr}$), such as MATH (Hendrycks et al., 2021b) and Numina-Math (Beeching et al., 2024). We generate additional math problems applying the similar technique as Ge et al. (2024) and define it as Persona-math ($\mathcal{D}_{syn\_mr}$).
Applying Templates for Answer Space and Reasoning Diversity.
General purpose reasoning benchmarks are often divided into two categories: (a) Multiple Choice Questions (Hendrycks et al., 2021a; Wang et al., 2024) and (b) Open-Ended Questions (Zhong et al., 2023). Recent works have ignored these variations in the answer space for consistent reward design across all tasks which are often predominantly math tasks (Hu et al., 2025; Aggarwal & Welleck, 2025; Luo et al., 2025). We hypothesize that each question type elicits different thinking patterns, leading to diverse reasoning trajectories in the model. Training on different question types will enhance the model’s ability to generalize by exposing it to diverse answer formats, thereby fostering different reasoning pathways.
Therefore, to observe the effect of question type in rl training, we synthesize $\mathcal{D}_{gpr}$ using two templates: $\mathcal{T}_{MCQ}$ Multiple Choice Questions (mcq), and $\mathcal{T}_{Open}$ Open-Ended questions. We convert the mcq datasets (mmlu) to open-ended by removing the options from the questions.
Additionally, some mcq questions are incomplete without options (e.g., Which of the following ways we can file taxes?). We discard such questions to avoid confusion during answer generation. Finally, our general purpose reasoning data, $\mathcal{D}_{gpr}$, can be represented as:
@@
To ensure high-quality training data, we apply a series of filtering and formatting steps, $\mathcal{H}$, to remove samples that are infeasible to evaluate with a simple rule-based reward function. Specifically, for $\mathcal{D}_{mcq}$, we check whether the correct answer appears within the question text itself. Given a question-answer pair $(q,a^{*})$ with answer choices $\{a_{1},a_{2},\dots,a_{n}\}$, we discard a sample if $a^{*}\notin\{a_{1},a_{2},\dots,a_{n}\}$.
For $\mathcal{D}_{open}$, such as those in the Natural Reasoning dataset, we discard samples that are challenging to evaluate with a rule-based reward function. Formally, we retain samples where $|w(a^{*})|\leq 10$; $w(a^{*})$ represents the number of words in the answer $a^{*}$.
Lastly, for the mathematical reasoning corpus, $\mathcal{D}_{mr}$, we remove entries that lack an associated answer, ensuring that all retained questions $q$ have a valid response $a^{*}$, i.e., we discard samples where $a^{*}=\emptyset$.
$\mathcal{D}^{\prime}=\mathcal{H}(\mathcal{D})=\left\{(q,a^{*},\{a_{1},\dots,a_{n}\})\in\mathcal{D}\;\middle|\;\begin{array}[]{l}a^{*}\in\{a_{1},\dots,a_{n}\}\quad(\mathcal{D}_{mcq})\\ |w(a^{*})|\leq 10\quad(\mathcal{D}_{open})\\ a^{*}\neq\emptyset\quad(\mathcal{D}_{mr})\end{array}\right.$

### v3

@@ Abstract
Prior work has successfully applied Reinforcement Learning (rl) to mathematical reasoning—where rules and correctness are well-defined. Yet, generalizing these methods to broader reasoning domains remains challenging due to limited data and the lack of verifiable rewards for unstructured domains. In this work, we propose Nemotron-CrossThink, a framework that systematically incorporates multi-domain corpora into rl training to improve generalization across diverse reasoning tasks. Nemotron-CrossThink addresses key challenges by (1) combining data from varied sources; (2) applying structured templates to control answer-space complexity; (3) filtering for verifiable answers; and (4) optimizing data blending strategies to utilize multi-source data effectively. This enables scalable and verifiable reward modeling beyond math and demonstrates improved accuracies on both math (math-500: +30.1%, amc23: +27.5%) and non-math reasoning benchmarks (mmlu-pro: +12.8%, gpqa-diamond: +11.3%, agieval: +15.1%, supergpqa: +3.8%). Moreover, Nemotron-CrossThink exhibits significantly improved response efficiency—using 28% fewer tokens for correct answers—highlighting more focused and effective reasoning. Through Nemotron-CrossThink, we demonstrate that integrating multi-domain, multi-format data in rl leads to more accurate, efficient, and generalizable llms. All of our datasets are available on HuggingFace.
Nemotron-CrossThink: Scaling Self-Learning beyond Math Reasoning
Syeda Nahida Akter2thanks: Work done during internship at NVIDIA,   Shrimai Prabhumoye1,3,   Matvei Novikov1,   Seungju Han1, Ying Lin1,   Evelina Bakhturina1,   Eric Nyberg2,   Yejin Choi1,   Mostofa Patwary1, Mohammad Shoeybi1,   Bryan Catanzaro1 NVIDIA1, Carnegie Mellon University2, Boston University3 sakter@andrew.cmu.edu,  sprabhumoye@nvidia.com
@@ 1 Introduction
Large Language Models (llms) have demonstrated remarkable reasoning abilities across a wide range of tasks, with Reinforcement Learning (rl) playing a key role in refining their deep thinking abilities (Hu et al., 2025a; Aggarwal and Welleck, 2025; Luo et al., 2025; DeepSeek-AI, 2025; Qin et al., 2024; Huang et al., 2025; Team, 2025c). Recent advances in rl have been particularly successful in mathematical reasoning and coding, where well-defined rules and verifiable correctness enable effective reward modeling. Yet, extending these techniques to broader reasoning domains poses significant challenges, such as—limited training data for rl due to the difficulty of defining verifiable rewards, and ensuring generalization across diverse tasks.
@@
Recent works (Hu et al., 2025a; Luo et al., 2025; Cui et al., 2025) have shown ways to diversify rl training corpora by collecting datasets from multiple sources. However, they do not evaluate the relative importance of each source for reasoning or explore optimal data-blending strategies to maximize performance. Furthermore, prior research has largely focused on math reasoning, overlooking the role of non-math reasoning domains in rl training for generalization in out-of-distribution domains. Reasoning process varies across domains and question types. For instance, math problem-solving follows a rule-based, structured, and symbolic approach (Dehaene, 2011), whereas reasoning in fields such as law, physics, social sciences, and history often relies on narrative structures, contextual knowledge, and heuristic strategies. Moreover, different question formats require distinct cognitive approaches—open-ended questions demand the generation of novel responses, while multiple-choice questions (mcq) can often be solved by evaluating the given options and selecting the most appropriate answer. Incorporating a diverse range of reasoning domains and question types into rl-based self-learning can enhance the broad reasoning ability of llms by exposing them to varied cognitive strategies and knowledge structures.
In this work, we propose Nemotron-CrossThink, a systematic way to incorporate multi-domain corpora for rl training that results in better generalization across a wide variety of tasks. As outlined in Figure 2, Nemotron-CrossThink comprises of phases that—(a) curate data from diverse sources, including synthetic data from web texts and open-source question-answer (QA) pairs on STEM, humanities, law, and social sciences (b) apply templates (mcq/Open-Ended) to limit the answer-space (d) prepare blends to combine multi-domain data efficiently and (e) employ self-learning with rl to refine reasoning capability in diverse domains.
@@
We evaluate Nemotron-CrossThink along three axes: (1) the effectiveness of data blending strategies in self-learning (2) whether the blending impact amplifies by training with more complex data samples (3) the influence of question-answer templates on downstream accuracies. Nemotron-CrossThink demonstrates that the integration of multi-domain data with varied question formats in rl boosts llm’s reasoning across diverse domains (Figure 1). Notably, models trained with Nemotron-CrossThink not only achieve higher accuracy but also exhibit dynamic response strategies—generating concise answers for general-purpose questions and more detailed responses for math problems—thereby reducing inference cost while preserving task-specific rigor. In addition, our approach addresses the challenge of designing scalable verifiable reward for non-deterministic domains by employing different templates on the curated data to limit the nuances in the answer space diversity. Furthermore, we explore a simple yet effective filtering approach to rank general purpose reasoning (gpr) data based on complexity and show that training with harder samples further amplifies the impact of rl across all domains.
In summary, our contributions are as follows:
@@
  • We introduce Nemotron-CrossThink, a novel framework for incorporating multi-domain corpora into rl training, enhancing generalization across diverse reasoning tasks with substantial gains on math (math-500: +30.1%, amc23: +27.5%) and non-math (mmlu-pro: +12.8%, gpqa-diamond: +11.3%, agieval: +15.1%, and supergpqa: +3.8%) benchmarks.
@@
  • We demonstrate that applying question/answer templates to constrain output diversity leads to more stable reward modeling. Specifically, using a unified open-ended question format improves performance by 1.21% over mixed-format questions, while short-form answer templates outperform long-form ones by 1.20%.
@@
  • We show that math-only training is insufficient—blending multi-domain data in rl boosts average reasoning accuracy by 1.61% over math-only data and improves response efficiency by reducing token usage by 28%.
@@
  • We release 287.4K high-quality multi-domain data curated for verifiable reward modeling to support future research [^1] https://huggingface.co/datasets/nvidia/Nemotron-CrossThink.
Applying Nemotron-CrossThink on different blends yields substantial improvement over base model (+8.55%-13.36% on average) across seven diverse gpr and math benchmarks. The most effective blend—2:1 ratio of gpr to math data—achieves the highest average accuracy with a 13.36% gain over baseline (Figure 1). Overall, these findings illustrate that thoughtful choices in data blending, scaling, formatting, and filtering are critical to the success of rl with language models. We hope that Nemotron-CrossThink serves as a practical and extensible framework for leveraging multi-domain data to train more capable, reliable, and generalizable models under the rl paradigm.
@@ 2 Nemotron-CrossThink: Scaling Self-Learning Beyond Math
Figure 2: Nemotron-CrossThink. We (a) curate QA pairs from from CommonCrawl and open-source datasets, categorized into general-purpose reasoning ($\mathcal{D}_{gpr}$) and mathematical reasoning ($\mathcal{D}_{mr}$); (b) apply structured templates to convert data into mcq and open-ended formats, promoting diverse reasoning trajectories; (c) filter out unverifiable or ill-formatted responses; (d) deploy rl using Group Relative Policy Optimization (grpo). The final reward is used to update the policy, iteratively improving the model’s reasoning capabilities across diverse domains.
While mathematical reasoning benefits from clean, verifiable datasets, extending rl to general-purpose reasoning is challenging due to the lack of structured, high-quality supervision. To address this, we leverage web documents and open-source QA benchmarks to collect general-purpose reasoning (gpr) data. However, combining structured and unstructured domains introduces noise and ambiguity—particularly in open-ended formats—making it difficult to apply rule-based reward reliably. To mitigate this, we apply task-specific templates to unify formats, limiting answer space variability and enabling effective verifiable reward signals. Next, we apply a lightweight data filtering to discard unverifiable examples for stable and interpretable rl training. Finally, we explore optimal data blending strategies to investigate how the inclusion of general-purpose reasoning data complements mathematical reasoning, ultimately leading to broader and more adaptive generalization in llms.
@@
As shown in Table 1, we start with curating datasets from multiple sources to ensure diversity in the training data. Our training data $\mathcal{D}$ comprises of:
@@
Here, $\mathcal{D}_{syn}\rightarrow$ synthetically generated from CommonCrawl (CC) (Gao et al., 2020) and $\mathcal{D}_{os}\rightarrow$ open-source QA datasets. Each source of data further consists of QA pairs related to gpr and math:
$\mathcal{D}_{\text{syn}}$ $\rightarrow\mathcal{D}_{\text{syn\_gpr}}\cup\mathcal{D}_{\text{syn\_mr}}$
$\mathcal{D}_{\text{os}}$ $\rightarrow\mathcal{D}_{\text{os\_gpr}}\cup\mathcal{D}_{\text{os\_mr}}$
  • General Purpose Reasoning, $\mathcal{D}_{gpr}$: We collect open source QA datasets ($\mathcal{D}_{os\_gpr}$)—Natural Reasoning (Yuan et al., 2025) and mmlu [Train] (Hendrycks et al., 2021a) that span domains including STEM, Economics, Social Sciences, and more. To enhance diversity, we further synthesize QA pairs from CC documents called Nemotron-CrossThink-qa ($\mathcal{D}_{syn\_gpr}$).
  • Mathematical Reasoning, $\mathcal{D}_{mr}$: We combine open-source math datasets ($\mathcal{D}_{os\_mr}$): MATH (Hendrycks et al., 2021b) and Numina-Math (Beeching et al., 2024). We generate additional math problems defined as Nemotron-CrossThink-math ($\mathcal{D}_{syn\_mr}$) to augment reasoning diversity.
Applying Templates for Answer Space and Reasoning Diversity.
General purpose reasoning benchmarks are often divided into two categories: (a) Multiple Choice Questions (Hendrycks et al., 2021a; Wang et al., 2024) and (b) Open-Ended Questions (Zhong et al., 2023). Prior works overlooked these variations in the answer space for consistent reward design for tasks which are predominantly math (Hu et al., 2025a; Aggarwal and Welleck, 2025; Luo et al., 2025). We hypothesize that each question type elicits different thinking patterns, leading to diverse reasoning trajectories in the model. Therefore, we synthesize $\mathcal{D}_{gpr}$ using two templates: $\mathcal{T}_{MCQ}$ - Multiple Choice Questions (mcq), and $\mathcal{T}_{Open}$ - Open-Ended questions. We convert the mcq datasets (mmlu) to open-ended by removing the options from the questions.
@@
Table 1: Training data distribution by source and type. oe=Open-Ended; gpr=General-Purpose Reasoning; mr=Math Reasoning.
Additionally, some mcq questions are incomplete without options (e.g., Which of the following ways we can file taxes?). We discard them to avoid confusion during answer generation. Finally,
@@
To obtain high-quality data, we apply a series of filtering and formatting steps, $\mathcal{H}$, to remove samples that are infeasible to evaluate with rule-based reward. Specifically, for $\mathcal{D}_{mcq}$, we check whether the correct answer appears within the question text itself. Given a question-answer pair $(q,a^{*})$ with answer choices $\{a_{1},a_{2},\dots,a_{n}\}$, we discard a sample if $a^{*}\notin\{a_{1},a_{2},\dots,a_{n}\}$. For $\mathcal{D}_{open}$, we discard samples that are challenging to evaluate with a rule-based reward function. Formally, we retain samples where $|w(a^{*})|\leq 10$; $w(a^{*})$ represents the number of words in the answer $a^{*}$.
Lastly, for $\mathcal{D}_{mr}$, we remove entries that lack an associated answer, ensuring that all retained questions $q$ have a valid response $a^{*}$, i.e., we discard samples where $a^{*}=\emptyset$.
$\mathcal{D}^{\prime}=\mathcal{H}(\mathcal{D})=\left\{(q,a^{*},\{a_{1},\dots,a_{n}\})\in\mathcal{D}\right$.